OpenStudy (anonymous):

For the limit lim x → 0 e2x − 1 x = 2 illustrate the definition by finding the largest possible values of δ that correspond to ε = 0.4 and ε = 0.3

OpenStudy (anonymous):

\[e ^{2x}-1/x=2\]

OpenStudy (ash2326):

\[\lim_{x \to 0} \frac{e^{2x}-1}{x}=2\]

OpenStudy (anonymous):

yeah e power 2x-1/=2

OpenStudy (anonymous):

any idea?